OpenStudy (anonymous):

a particle is placed at a point O on the edge of a horizontal table at height 2a above a horizontal floor. the point on the floor vertically below O is A. the particle is projected horizontally away from the table and lands on the floor at point B, where AB=a. show that the speed of projection is (1/2)(ga)^1/2. find the magnitude and direction of the velocity of the particle at the instant when it is vertically above the mid-point of AB. at time t after the projection the angle PAB is θ. show that tan θ= (1/(ga)^1/2)((4a/t)-gt)

OpenStudy (anonymous):

vertical 2a = 1/2 *gt^2 t=sqrt(4a/g) horizontal a = V*t V=a/t = a/(sqrt(4a/g) = 1/2 * sqrt(ag)

OpenStudy (anonymous):

now that you have initial horizontal velocity, try the second part... horiontal a/2 = 1/2 * sqrt(ag)*t t = sqrt(a)/sqrt(g) vertical Vf = g*t Vf = g* sqrt(a)/sqrt(g) = sqrt(ag)

OpenStudy (anonymous):

find the magnitude of velocity using those two components...

OpenStudy (anonymous):

ok i will

OpenStudy (anonymous):

let me now if you're having trouble with the last bit, the angle...